# Extending edge colorings of distance-3 matchings in the Cartesian product of graphs

### v4

@@ 1 Introduction
Definition 1.1.
We call a precoloring of the edges in $E_{0}$ extendable if there is a proper edge coloring with $\chi^{\prime}(G)$ colors such that the color of the edges in $E_{0}$ are the prescribed colors. Such a coloring is called the extension of the precoloring.
@@
The Cartesian product $G_{1}\square G_{2}$ of graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph such that the vertex set of $G_{1}\square G_{2}$ is the Cartesian product $V(G_{1})\times V(G_{2})$ and two vertices $(u_{1},u_{2})$ and $(v_{1},v_{2})$ are adjacent in $G_{1}\square G_{2}$ if and only if either
@@
Notation 1.3.
$G^{d}$ denotes the $d$th power of the Cartesian product of $G$ with itself.
Notation 1.4.
One of the basic concepts in the paper is as follows.
Definition 1.5.
(The brick-neighborhoods of an edge in the Cartesian product) Let $yz\in E(G)$ be an edge. Then take a path $xyzw$ in $G$ and an edge $pq\in E(H)$. Then a brick-neighborhood of the edge $(y,q)(z,q)$ is the Cartesian product of the path $xyzw$ and the edge $pq$ ($xyzw\square pq$). Symmetrically, we can define brick-neighborhoods starting with an edge in $H$ (i.e., we take a 3-path in $H$ and an edge in $G$).
Remark 1.6.
Definition 1.7.
Problem 1.8.
@@
Casselgren, Granholm, and Petros [2] proved
@@
In the paper [2], the authors raised the conjecture as follows.
@@
We recall that the direct product of a Class 1 graph with any other graph is also a Class 1 graph. Therefore, in the case of a product with multiple factors, if at least one of the factors is Class 1, then the entire product is Class 1 as well.
@@
Let $k$ be an integer with $k\geq 2$, and let $G_{i}$ be Class 1 graphs for each $i\in\{1,\dots,k\}$, with maximum degree $\Delta(G_{i})$, respectively. Suppose that some edges of the Cartesian product $G_{1}\square G_{2}\square\cdots\square G_{k}$ are precolored using at most $\sum_{i=1}^{k}\Delta(G_{i})$ colors, ensuring that the distance between any two precolored edges is at least three. If $\sum_{i=1}^{k}\Delta(G_{i})$ is even and satisfies $2\Delta(G_{i})<\sum_{i=1}^{k}\Delta(G_{i})\forall i\in\{1,\dots,k\}$, then this precoloring is extendable.
@@
Remark 1.17.
Theorem 1.15 solves Problem 1.8 in the case where $\Delta(G)=\Delta(H)$ and also Conjecture 1.12, as the degree conditions are clearly satisfied in both cases.
In the second section, we also study the problem in the case of the Cartesian product of a bipartite graph and an edge. Galvin’s theorem [7] is used in the proof of this statement, so the proof does not work for the Cartesian product of a Class 1 graph and an edge. However, if the List Coloring Conjecture holds for Class 1 graphs, then the proof automatically extends to this case as well. The final section presents some related results on Cartesian products with odd cycles. The following definition will be useful in the proofs.
Definition 1.18.
(Canonical coloring of the Cartesian product of graphs) Let us fix a proper coloring of each Class 1 graph $G_{i}$ for $i\in\{1,\dots,k\}$ using $\Delta(G_{i})$ colors, respectively, resulting in a total of $\sum_{i=1}^{k}\Delta(G_{i})$ distinct colors. We define the canonical coloring of the graph $G=G_{1}\square G_{2}\square\cdots\square G_{k}$ as follows. If an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, we assign it the same color as the edge $x_{j}y_{j}$ in $G_{j}$.
@@ 2 Admissible edge-colorings of the Cartesian
Given $k\geq 2$ disjoint sets $H_{1},\dots,H_{k}$ and their union $H=H_{1}\cup\dots\cup H_{k}$ where $|H|$ is even and satisfies $2|H_{i}|<|H|$ for all $i\in\{1,\dots,k\}$. We can pair off the elements of $H$ so that no two elements in any pair belong to the same set $H_{i}$.
@@
In the proof, we will show that a single factor can be made regular. This procedure can then be applied individually to each factor, thereby proving the statement.
We describe a possible procedure for adding these extra vertices and edges to a graph $H$. Starting with $H$, whenever a vertex has a degree less than $\Delta(H)$, we duplicate the entire graph and connect each such vertex to its counterpart in the copy. Duplicating $H$ ensures that any vertex of degree less than $\Delta(H)$ gains exactly one new neighbor, increasing its degree without exceeding $\Delta(H)$. This operation increases the minimum degree without raising the chromatic index. We repeat this process until the graph becomes regular.
It is clear that the graph remains Class 1 at every step, since we can color both copies identically. Furthermore, since new edges are only added between vertices whose degrees are still below $\Delta(H)$, we can assign proper colors to these new edges without increasing the total number of colors used.
Let $G$ and $H$ be Class 1 graphs, and let $H^{*}$ denote the regular extension of $H$ obtained by this process. If two vertices had distance at least 3 in the graph $G\square H$, then their distance will also be at least 3 in the graph $G\square H^{*}$.
Distance one is clearly impossible, since no new edges are added within the $G\square H$ subgraph. Distance two is also not possible, as this would require one of the following two configurations:
  • In the first case, the three vertices are $(u,v)$, $(u,w)$, and $(u,z)$, where $w$ must necessarily be a new vertex. Otherwise, this would already be a path of length two in $G\square H$, contradicting our assumption. However, based on the construction, each new vertex has at most one neighbor in the original graph, making this configuration impossible.
  • In the second case, the three vertices are $(u,v)$, $(u,w)$, and $(y,w)$. But in this case neither $u$ nor $w$ can be a new vertices, again contradicting the rules of the construction. Hence, this configuration is also impossible.
@@
We can color the components of the graph one by one, so we may assume that the graph is connected. Let $G=G_{1}\square\dots\square G_{k}$ be used in the following. Consider a canonical coloring of the graph $G$. This is a proper coloring with $\Delta(G)$ colors, but is usually not an admissible coloring. Now, we modify the coloring step by step, preserving the number of colors and ensuring that the coloring is proper. If a precolored edge has the correct color, then we will not change it. If the prescribed color of a precolored edge is not its color in the canonical coloring, then we will correct it by some local operations. So, we finally get an admissible coloring.
@@
STEP 1: Suppose that an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, and its prescribed color appears in the coloring of the graph $G_{l}$ for some $l\neq j$. Since $G_{l}$ is regular, there must be an edge $(x_{l},z)\in E(G_{l})$ that got this prescribed color in the canonical coloring. In this case, we perform a rotation on the square $x_{j}y_{j}\square x_{l}z$. This rotation affects only the internal edges of a brick-neighborhood of the precolored edge, ensuring that the brick-neighborhoods of the other precolored edges remain unchanged (see Remark 1.16).
STEP 2: Suppose that an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, and its prescribed color appears in the coloring of the graph $G_{j}$, but the color of this edge in the canonical coloring is different from the prescribed color. Then there are edges $sx_{j},y_{j}t\in E(G_{j})$, which have this color in the canonical coloring. Now, take an arbitrary edge $(x_{i},w)$ of the graph $G_{i}$ for some $i\neq j$ and perform a rotation sequentially on the squares $sx_{j}\square x_{i}w$, $y_{j}t\square x_{i}w$ and finally on $x_{j}y_{j}\square x_{i}w$.
@@
In this case, the coloring changes on the external edges, as well. Therefore, we aim to select one of the brick-neighborhoods of each of these type precolored edges in such a way that they are pairwise edge-disjoint. If we can do this, then we can achieve an admissible coloring by applying the second type of step above to the brick-neighborhoods chosen in this way.
The degrees of the graphs $G_{1},\dots,G_{k}$ (and consequently the colors used) satisfy the conditions of Lemma 2.1 Therefore, the colors used for the coloring $G_{1},\dots,G_{k}$ can be paired so that no two colors in any pair originate from the same graph $G_{i}$. For a precolored edge $yz$, we select the brick-neighborhood $B=xyzw\square pq$ such that the color of $xy$ and $zw$ is the prescribed color of $yz$ and the color of $pq$ is the pair of this color. Let $f(B)$ denote this pair of colors, i.e., the colors assigned to the edges of the end squares $xy\square pq$ or $zw\square pq$. For any two chosen brick-neighborhoods $B_{1}$ and $B_{2}$ (corresponding to precolored edges), sets $f(B_{1})$ and $f(B_{2})$ are either identical or disjoint.
Notice that any two selected brick-neighborhoods are edge-disjoint. Suppose, for contradiction, that two brick-neighborhoods share an external edge (by Remark 1.16) In that case, they must have two end squares sharing exactly one edge, leading to two color pairs that share exactly one color, which contradicts our pairing condition. This completes the proof of Theorem 1.15. ∎
@@ 3 Admissible edge-colorings of the Cartesian
This is the strongest possible statement. The graph $C_{2k+1}\square C_{2l+1}$ has an odd number of vertices, which means that it does not contain perfect matching. Therefore, it cannot be 4-edge-colorable, as that would require its edges to be divided into four perfect matchings. It is easy to give a non-extendable precoloring of the graph $C_{2k+1}\square C_{2l+1}$ where the minimum distance between precolored edges is at least two (see Remark 1.13 or [2] ).
@@
[…]
@@
Afterward, the edges of types $(i,2l)(i,0)$ and $(2k,j)(0,j)$ are colored with the fifth color. (This will not be a proper coloring, but aside from the edges of the square $(2k,2l)(2k,0)(0,0)(0,2l)$, it is proper.) If an edge $(i_{1},j_{1})(i_{2},j_{2})$, where $\{i_{1},j_{1},i_{2},j_{2}\}\cap\{0,2k,2l\}=\emptyset$, has the fifth color prescribed, it can be colored with this color and no further changes are needed. If an edge $(i_{1},j_{1})(i_{2},j_{2})$ has exactly one of $i_{1},j_{1},i_{2},j_{2}$ from the set $\{0,2k,2l\}$ and has the fifth color prescribed, it is adjacent to exactly one edge of this color. In this chase, we can swap the color of these two edges. Similarly, if an edge of type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ has a prescription such that it has exactly one adjacent edge of this color, then we can swap the color of these two edges. If an edge of type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ has a prescription such that it has two adjacent edges of this color, then they define a square, and we can rotate on this square. If there is a prescription of the fifth color on an edge of type $(i,*)(j,*)$ or $(*,i)(*,j)$ (the symbol $*\in\{0,2k,2l\}$ denotes the same value in the two vertices), then there is a square with two sides colored with the fifth color; we can rotate on this square.
@@
Now, consider the squares $A=(2k-1,0)(2k,0)(2k,1)(2k-1,1)$ and $B=(0,0)(0,1)(1,1)(1,0)$. If there exists a color other than $5$ that is not used in the coloring of either $A$ or $B$, then we can assign that color to the edge $(2k,0)(0,0)$. In fact, we only need to check the colors of the edges incident to the vertices $(2k,0)$ or $(0,0)$.
In the canonical coloring, the colors $2$ and $3$ appear in the coloring of the square $A$, while the colors $1$ and $3$ appear in the coloring of the square $B$. If these colors remain unchanged, we can color the edge $(2k,0)(0,0)$ with the color $4$. However, this situation changes only if at least one of the edges $(2k-1,0)(2k-1,1)$, $(2k-1,1)(2k,1)$, $(2k,1)(0,1)$, $(0,1)(1,1)$ and $(1,1)(1,0)$ is precolored.
If the edges $(2k-1,0)(2k,0)$ and $(2k,0)(2k,1)$ or $(0,0)(0,1)$ and $(0,0)(1,0)$ are external edges in the brick-neighborhood of a precolored edge, this will not affect the overall coloring. During rotations, the colors of these edges may change, but they will still have the same two colors, just in a different order.
@@
  • The edge $(2k-1,1)(2k,1)$ is prescribed to be color $1$, and we performed the rotations in the brick-neighborhood containing $A$. As a result, the square $A$ will be colored with the colors $1$ and $2$, while the square $B$ remains unchanged, colored with the colors $1$ and $3$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $4$.
  • The edge $(2k-1,0)(2k-1,1)$ is prescribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $A$, with no changes occurring in the square $B$. As a result, the square $A$ will be colored with colors $3$ and $4$, while the square $B$ remains unchanged, colored with the colors $1$ and $3$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $2$.
  • The edge $(2k-1,0)(2k-1,1)$ is precribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $A$. Similarly, the edge $(1,0)(1,1)$ is prescribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $B$. As a result, both square $A$ and $B$ will be colored with the colors $3$ and $4$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $2$.

### v5

@@ 1 Introduction
We call a precoloring of some edge set $E_{0}$ extendable if there is a proper edge coloring with $\chi^{\prime}(G)$ colors such that the color of the edges in $E_{0}$ are the prescribed colors. Such a coloring is called an extension of the precoloring.
@@
The Cartesian product $G_{1}\square G_{2}$ of graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph whose vertex set is the Cartesian product $V(G_{1})\times V(G_{2})$ and two vertices $(u_{1},u_{2})$ and $(v_{1},v_{2})$ are adjacent in $G_{1}\square G_{2}$ if and only if either
@@
Notation 1.2.
$G^{d}$ denotes the $d$-th power of the Cartesian product of $G$ with itself.
Problem 1.3.
@@
Casselgren, Granholm, and Petros [2] proved the following.
@@
In the paper [2], the authors raised a conjecture as follows.
@@
We recall that the Cartesian product of a Class 1 graph with any other graph is also a Class 1 graph. Therefore, in the case of a product with multiple factors, if at least one of the factors is Class 1, then the entire product is Class 1 as well.
@@
Let $k$ be an integer with $k\geq 2$, and let $G_{i}$ be Class 1 graphs for each $i\in\{1,\dots,k\}$, with maximum degree $\Delta(G_{i})$, respectively. Suppose that some edges of the Cartesian product $G_{1}\square G_{2}\square\cdots\square G_{k}$ are precolored using at most $\sum_{i=1}^{k}\Delta(G_{i})$ colors, ensuring that the distance between any two precolored edges is at least three. If $\sum_{i=1}^{k}\Delta(G_{i})$ is even and satisfies $2\Delta(G_{i})<\sum_{j=1}^{k}\Delta(G_{j})\quad\forall i\in\{1,\dots,k\}$, then this precoloring is extendable.
@@
Theorem 1.10 solves Problem 1.3 in the case where $\Delta(G)=\Delta(H)$ and also Conjecture 1.7, as the degree conditions are clearly satisfied in both cases.
In the second section, we also study the problem in the case of the Cartesian product of a bipartite graph and an edge. Galvin’s theorem [7] is used in the proof of this statement, so the proof does not work for the Cartesian product of a Class 1 graph and an edge. However, if the List Coloring Conjecture holds for Class 1 graphs, then the proof automatically extends to this case as well. The final section presents some related results on Cartesian products with odd cycles.
@@ 2 Admissible edge-colorings of the Cartesian
Notation 2.1.
One of the basic concepts in the paper is as follows.
Definition 2.2.
(The brick-neighborhoods of an edge in the Cartesian product)
Let $yz\in E(G)$ be an edge. Then take a path xyzwxyzw in $G$ and an edge $pq\in E(H)$. Then a brick-neighborhood of the edge $(y,q)(z,q)$ is the Cartesian product of the path xyzwxyzw and the edge pqpq ($xyzw\square pq$). Symmetrically, we can define brick-neighborhoods starting with an edge in $H$ (i.e., we take a 3-path in $H$ and an edge in $G$).
Remark 2.3.
Remark 2.4.
The following definition will be also useful in the proofs.
Definition 2.5.
(Canonical coloring of the Cartesian product of graphs)
Let us fix a proper coloring of each Class 1 graph $G_{i}$ for $i\in\{1,\dots,k\}$ using $\Delta(G_{i})$ colors, respectively, resulting in a total of $\sum_{i=1}^{k}\Delta(G_{i})$ distinct colors. We define the canonical coloring of the graph $G=G_{1}\square G_{2}\square\cdots\square G_{k}$ as follows. If an edge $(x_{1},\dots,x_{k})(y_{1},\dots,y_{k})\in E(G)$ connects two vertices that differ in the $j$th coordinate, we assign it the same color as the edge $x_{j}y_{j}$ in $G_{j}$.
@@
Given $k\geq 2$ disjoint sets $H_{1},\dots,H_{k}$ and their union $H=H_{1}\cup\dots\cup H_{k}$ where $|H|$ is even and satisfies $2|H_{i}|<|H|$ for all $i\in\{1,\dots,k\}$. The elements of $H$ can be partitioned into $|H|/2$ disjoint pairs such that no pair is a subset of any $H_{i}$.
@@
It suffices to show that a single factor can be made regular. This procedure can then be applied individually to each factor, thereby proving the statement.
We describe a possible procedure for adding these extra vertices and edges to a graph $H$. Starting with $H$, whenever a vertex has a degree less than $\Delta(H)$, we duplicate the entire graph and connect each such vertex to its counterpart in the copy. This ensures that any vertex of degree less than $\Delta(H)$ gains exactly one new neighbor, increasing its degree without exceeding $\Delta(H)$. This operation increases the minimum degree without raising the chromatic index. We repeat this process until the graph becomes regular. (The minimum degree increases in every step and the maximum degree is unchanged, so the process stops in $\Delta(H)-\delta(H)$ steps.)
We claim that the graph remains Class 1 at every step. $H$ has a $\Delta(H)$-edge-coloring and we can color both copies identically. Furthermore, new edges are only added between vertices whose degrees are still below $\Delta(H)$, so at each such vertex there is an unused color. We can assign these unused colors to these new edges without increasing the total number of colors used.
Now, let $G$ and $H$ be Class 1 graphs and let $H^{*}$ denote the regular extension of $H$ obtained by this process. We show that if the distance between two edges $e_{1},e_{2}$ in $G\square H$ is at least 3, it remains at least 3 in $G\square H^{*}$. Assume, for contradiction, that $d_{G\square H}(e_{1},e_{2})\geq 3$ but $d_{G\square H^{*}}(e_{1},e_{2})\leq 2$.
Since no new edges were added between original vertices, the distance cannot be 1. (And the distance cannot be 0, of course.) If the distance is 2, there must exist a vertex $(a,b)$ in $G\square H^{*}$ and vertices $(u,v)\in e_{1}$ and $(x,y)\in e_{2}$ such that $(u,v)\sim(a,b)$ and $(a,b)\sim(x,y)$.
  1. 1.
    Case 1: $(a,b)$ is an original vertex in $G\square H$. Since the construction only adds edges between original and new vertices (or between two new vertices), the adjacency relations between original vertices are unchanged. This implies $d_{G\square H}((u,v),(x,y))=2$, contradicting $d_{G\square H}(e_{1},e_{2})\geq 3$.
  2. 2.
    Case 2: $(a,b)$ is a new vertex. By the definition of the Cartesian product, for the adjacency $(u,v)\sim(a,b)$ where $(u,v)$ is original and $(a,b)$ is new, we must have $u=a$ and $v\sim b$ in $H^{*}$. Similarly, for $(x,y)\sim(a,b)$, we must have $x=a$ and $y\sim b$ in $H^{*}$. In our regularization procedure, whenever we duplicate a graph, we only add edges between a vertex and its direct copy (or between two new vertices). By induction, any new vertex $b\in V(H^{*})\setminus V(H)$ is connected to at most one vertex in the original vertex set $V(H)$, which is its unique ”preimage” (the vertex from which it was originally derived). Since $v\sim b$ and $y\sim b$ in $H^{*}$, and $v,y\in V(H)$ while $b\notin V(H)$, the construction implies that $v$ and $y$ both must be the unique preimage of $b$ in $V(H)$. It follows that $u=x$ and $v=y$. Thus, $(u,v)=(x,y)$, which means that $e_{1}$ and $e_{2}$ share a vertex, so their distance is 0, a contradiction.
Since all cases lead to a contradiction, our assumption was false: the distance cannot decrease below 3 in the regularized version. ∎
@@
Since the components of the graph can be colored independently, we may assume that $G=G_{1}\square\cdots\square G_{k}$ is connected. Consider a canonical coloring of the graph $G$. This is a proper coloring with $\Delta(G)$ colors, though it is not necessarily an admissible coloring. Now, we modify the coloring step by step, preserving the number of colors and ensuring that the coloring is proper. If a precolored edge has the correct color, then we will not change it. If the prescribed color of a precolored edge is not its color in the canonical coloring, then we will correct it by some local operations. So, we finally get an admissible coloring.
@@
Notation 2.8.
For a vertex $v\in V(G)$, we denote by $v^{(j:x)}$ the vertex identical to $v$ except in the $j$-th coordinate, where its value is $x$. Similarly, $v^{(i:x,j:y)}$ denotes the vertex where the $i$-th and $j$-th coordinates are replaced by $x$ and $y$, respectively.
From now on (unless otherwise stated), let the coordinates of a vertex $v$ be $v_{1},\dots,v_{k}$.
@@
STEP 1: Consider an edge $e=v^{(j:x)}v^{(j:y)}\in E(G)$ whose prescribed color appears in the coloring of $G_{l}$ for some $l\neq j$. Since $G_{l}$ is regular, there exists an edge $v_{l}z\in E(G_{l})$ colored with this color. By performing a rotation on the square with vertices $\{v^{(j:x)},v^{(j:y)},v^{(j:x,l:z)},v^{(j:y,l:z)}\}$, the edge $e$ receives its prescribed color. This operation modifies only the internal edges of the brick-neighborhood of $e$, ensuring that the coloring of the brick-neighborhoods of all other precolored edges remains unchanged (see Remark 2.4).
STEP 2: Consider an edge $e=v^{(j:x)}v^{(j:y)}\in E(G)$ whose prescribed color appears in the coloring of $G_{j}$, but differs from the color assigned to $e$ in the canonical coloring. Since $G_{j}$ is regular, there must exist edges $sx,yt\in E(G_{j})$, which have this prescribed color in the canonical coloring. Now, choose an arbitrary edge $v_{i}w$ of the graph $G_{i}$ for some $i\neq j$. To achieve the desired color on $e$, we perform a sequence of rotations (in the following order) on the squares defined by these vertex sets:
  1. 1.
    $\{v^{(j:s)},v^{(j:x)},v^{(j:s,i:w)},v^{(j:x,i:w)}\}$
  2. 2.
    $\{v^{(j:y)},v^{(j:t)},v^{(j:y,i:w)},v^{(j:t,i:w)}\}$
  3. 3.
    $\{v^{(j:x)},v^{(j:y)},v^{(j:x,i:w)},v^{(j:y,i:w)}\}$
@@
In this case, the coloring changes on the external edges as well. Therefore, we aim to select one of the brick-neighborhoods of each of these type precolored edges in such a way that they are pairwise edge-disjoint. If we can do this, then we can achieve an admissible coloring by applying STEP 2 to the brick-neighborhoods chosen in this way.
Let $H_{i}$ be the set of colors used in the coloring of $G_{i}$ for each $i=1,\dots,k$, and let $H=\bigcup H_{i}$. The sets $H_{i}$ satisfy the conditions of Lemma 2.6; therefore, there exists an involution $\varphi:H\to H$ (i.e. $\varphi(\varphi(x))=x$) such that if $x\in H_{i}$, then $\varphi(x)\notin H_{i}$. (This means that every pair of colors belongs to the colorings of two distinct graphs.)
In STEP 2, the index $i$ and the vertex $w$ can be freely chosen, provided that $i\neq j$ and $v_{i}w\in E(G_{i})$. We shall now provide a specific choice that guarantees that the brick-neighborhoods used are disjoint. For a precolored edge $e=v^{(j:x)}v^{(j:y)}\in E(G)$ with prescribed color $c$, let $sx,yt\in E(G_{j})$ be the edges with color $c$ in the canonical coloring. We select $i$ and $w$ so that the color of $v_{i}w\in E(G_{i})$ is $\varphi(c)$. We select the brick-neighborhood $B$ defined by the vertex set
$\{v^{(j:s)},v^{(j:x)},v^{(j:y)},v^{(j:t)},v^{(j:s,i:w)},v^{(j:x,i:w)},v^{(j:y,i:w)},v^{(j:t,i:w)}\}$
Let $f(B)=\{c;\varphi(c)\}$ denote the pair of colors associated with $B$. These are exactly the colors assigned to the edges of the squares involved in the first two rotations of STEP 2. For any two such brick-neighborhoods $B_{1}$ and $B_{2}$ (corresponding to precolored edges), sets $f(B_{1})$ and $f(B_{2})$ are either identical or disjoint. Specifically, if $c_{2}\in\{c_{1},\varphi(c_{1})\}$, then $\{c_{1},\varphi(c_{1})\}=\{c_{2},\varphi(c_{2})\}$; otherwise, $\{c_{1},\varphi(c_{1})\}\cap\{c_{2},\varphi(c_{2})\}=\emptyset$.
Finally, we show that any two selected brick-neighborhoods are edge-disjoint. Suppose, for contradiction, that two brick-neighborhoods share an external edge (by Remark 2.4). This implies that the two brick-neighborhoods contain a pair of (distinct) squares that share exactly one edge. Such squares can only be those involved in the first two rotations of STEP 2; therefore, the colors assigned to their edges in the canonical coloring belong to $f(B_{1})$ and $f(B_{2})$, respectively. As shown previously, these sets are either identical or disjoint. Since the two squares share an edge, which must have a unique color $c^{*}$, it follows that $c^{*}\in f(B_{1})\cap f(B_{2})$, which means $f(B_{1})=f(B_{2})$. However, in this case, the other two pairs of edges incident to the shared edge in each square would have the same color ($\varphi(c^{*})$). Because the squares are distinct and share only one edge, these are two pairs of distinct edges incident to the same vertex (see Figure 1). This means that two edges of the same color meet at a single vertex, which contradicts the fact that the coloring is proper in every step. This completes the proof of Theorem 1.10. ∎
$\varphi(c^{*})$$\varphi(c^{*})$$\varphi(c^{*})$$\varphi(c^{*})$$^{*}cc^{*}$$c^{*}$
Figure 1: Two squares from distinct brick-neighborhoods sharing an edge
@@ 3 Admissible edge-colorings of the Cartesian
This is the strongest possible statement. The graph $C_{2k+1}\square C_{2l+1}$ has an odd number of vertices, which means that it does not contain perfect matching. Therefore, it cannot be 4-edge-colorable, as that would require its edges to be divided into four perfect matchings.
@@
Label the vertices of the graph $C_{2k+1}\square C_{2l+1}$ with the labels $\{(i,j)\mid i=0\dots 2k;j=0\dots 2l\}$. The vertices $(x_{1},x_{2})$ and $(y_{1},y_{2})$ are joined by an edge if and only if $x_{1}=y_{1}$ and $|x_{2}-y_{2}|\equiv 1\ (\mathrm{mod}\ 2l+1)$ or $x_{2}=y_{2}$ and $|x_{1}-y_{1}|\equiv 1\ (\mathrm{mod}\ 2k+1)$.
@@
Afterward, the edges of types $(i,2l)(i,0)$ and $(2k,j)(0,j)$ are colored with the fifth color. (This will not be a proper coloring, but aside from the edges of the square $(2k,2l)(2k,0)(0,0)(0,2l)$, it is proper.) If an edge $(i_{1},j_{1})(i_{2},j_{2})$, where $\{i_{1},j_{1},i_{2},j_{2}\}\cap\{0,2k,2l\}=\emptyset$, has the fifth color prescribed, it can be colored with this color and no further changes are needed. If an edge $(i_{1},j_{1})(i_{2},j_{2})$ has exactly one of $i_{1},j_{1},i_{2},j_{2}$ from the set $\{0,2k,2l\}$ and has the fifth color prescribed, it is adjacent to exactly one edge of this color. In this chase, we can swap the color of these two edges. Similarly, if an edge of type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ has a prescription such that it has exactly one adjacent edge of this color, then we can swap the color of these two edges. If an edge of type $(i,2l)(i,0)$ or $(2k,j)(0,j)$ has a prescription such that it has two adjacent edges of this color, then they define a square, and we can rotate on this square. If there is a prescription of the fifth color on an edge of type (i,)(j,)(i,*)(j,*) or (,i)(,j)(*,i)(*,j) (the symbol $\in\{0,2k,2l\}$ denotes the same value in the two vertices), then there is a square with two sides colored with the fifth color; we can rotate on this square.
@@
Now, consider the squares A=(2k1,0)(2k,0)(2k,1)(2k1,1)A=(2k-1,0)(2k,0)(2k,1)(2k-1,1) and $B=(0,0)(0,1)(1,1)(1,0)$. If there exists a color other than $5$ that is not used in the coloring of either $A$ or $B$, then we can assign that color to the edge $(2k,0)(0,0)$. In fact, we only need to check the colors of the edges incident to the vertices $(2k,0)$ or $(0,0)$.
In the canonical coloring, the colors $2$ and $3$ appear in the coloring of the square $A$, while the colors $1$ and $3$ appear in the coloring of the square $B$. If these colors remain unchanged, we can color the edge $(2k,0)(0,0)$ with the color $4$. However, this situation changes only if at least one of the edges (2k1,0)(2k1,1)(2k-1,0)(2k-1,1), (2k1,1)(2k,1)(2k-1,1)(2k,1), $(2k,1)(0,1)$, $(0,1)(1,1)$ and $(1,1)(1,0)$ is precolored.
If the edges (2k1,0)(2k,0)(2k-1,0)(2k,0) and $(2k,0)(2k,1)$ or $(0,0)(0,1)$ and $(0,0)(1,0)$ are external edges in the brick-neighborhood of a precolored edge, this will not affect the overall coloring. During rotations, the colors of these edges may change, but they will still have the same two colors, just in a different order.
@@
  • The edge (2k1,1)(2k,1)(2k-1,1)(2k,1) is prescribed to be color $1$, and we performed the rotations in the brick-neighborhood containing $A$. As a result, the square $A$ will be colored with the colors $1$ and $2$, while the square $B$ remains unchanged, colored with the colors $1$ and $3$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $4$.
  • The edge (2k1,0)(2k1,1)(2k-1,0)(2k-1,1) is prescribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $A$, with no changes occurring in the square $B$. As a result, the square $A$ will be colored with colors $3$ and $4$, while the square $B$ remains unchanged, colored with the colors $1$ and $3$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $2$.
  • The edge (2k1,0)(2k1,1)(2k-1,0)(2k-1,1) is precribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $A$. Similarly, the edge $(1,0)(1,1)$ is prescribed to be color $4$, and we performed the rotations in the brick-neighborhood containing $B$. As a result, both square $A$ and $B$ will be colored with the colors $3$ and $4$. Therefore, the edge $(2k,0)(0,0)$ can be colored with $2$.